### v2

# A strengthening of the dimensional Brunn–Minkowski conjecture implies the (B)-conjecture

Abstract
In this note, we prove that if a sufficiently regular even log-concave measure satisfies a certain stronger form of the dimensional Brunn-Minkowski conjecture, then it also satisfies the (B)-conjecture. Furthermore, we show that hereditarily convex measures satisfy the aforementioned strengthened form, therefore providing an alternative proof of a recent result by Cordero-Erausquin and Eskenazis stating that a hereditarily convex measure satisfies both conjectures.
MSC 2020 Classification: 52A40, 52A20, 46N20, 47F10.
Keywords: convex bodies, Brunn–Minkowski inequality, B-inequality, Gardner–Zvavitch problem, log-concave measures, elliptic PDEs.
1 Introduction
A measure $\mu$ on $\mathbb{R}^{n}$ is said to be log-concave if it satisfies the inequality
$\mu((1-t)K+tL)\geq\mu(K)^{1-t}\mu(L)^{t}$ (1)
for every convex and compact sets $K,L\subset\mathbb{R}^{n}$ and every $t\in[0,1]$, where $K+L$ is the Minkowski sum of $K$ and $L$. Among the central open problems in the study of log-concave measures are the dimensional Brunn–Minkowski conjecture and the (B)-conjecture.
The dimensional Brunn–Minkowski conjecture asks whether the dimensional concavity of the Lebesgue measure extends to all log-concave measures under symmetry assumptions. More precisely, if a log-concave measure $\mu$ is even, that is, $\mu(A)=\mu(-A)$ for every $\mu$-measurable set $A\subset\mathbb{R}^{n}$, and if $K,L$ are origin-symmetric convex sets, then the conjecture asserts that
$\mu((1-t)K+tL)^{\frac{1}{n}}\geq(1-t)\mu(K)^{\frac{1}{n}}+t\mu(L)^{\frac{1}{n}}$ (dim-BM)
holds for every $t\in[0,1]$. Notice that (dim-BM) in the case of the Lebesgue measure reduces to the celebrated Brunn–Minkowski inequality (see, e.g., [16, Section 7]), which does not require any symmetry for $K$ and $L$. The dimensional Brunn–Minkowski conjecture was first stated by Gardner and Zvavitch [9] for the Gaussian measure, a case settled by Eskenazis and Moschidis [7], while Cordero-Erausquin and Rotem [6] proved (dim-BM) for a broader class of measures, including sufficiently regular rotationally invariant log-concave measures. Nevertheless, the conjecture remains open for general even log-concave measures. See [1, 5] for some recent developments and further references.
The (B)-conjecture, seemingly independent of (dim-BM), concerns the behaviour of suitable dilates of convex bodies with respect to a log-concave measure. It asks whether, for an even log-concave measure $\mu$ on $\mathbb{R}^{n}$, the function
$t\mapsto\mu(e^{t}K)\,,t\in\mathbb{R},$ (B)
is log-concave for every origin-symmetric convex $K\subset\mathbb{R}^{n}$. This problem was put forth by Banaszczyk and popularised by Latała [14]. For the case of the Gaussian measure, this question was affirmatively answered by Cordero-Erausquin, Fradelizi, and Maurey [4]. A previous proof by Brascamp and Lieb [3], which remained unnoticed for decades, recently resurfaced in [5]. Cordero-Erausquin and Rotem [6] later proved that the conjecture holds for the same general class of measures as (dim-BM) above. More recently, Cordero-Erausquin and Eskenazis [5] established the conjecture for so-called hereditarily convex measures (see later (4)), of which the measures treated in [6] are a subset.
In [5] it is shown that hereditarily convex measures satisfy both (dim-BM) and (B), extending earlier work of Kolesnikov and Livshyts [10], who proved that, for sufficiently regular log-concave measures, hereditary convexity implies (dim-BM). Although no direct implication between (dim-BM) and (B) is currently known, we show in this note that a suitable strengthening of (dim-BM) yields (B). Furthermore, we prove that this strengthening is implied by hereditary convexity, thereby providing an alternative route to some of the results in [5]. Our approach relies on a new method introduced in [1], combined with an application of the $L_{2}$ method developed by Kolesnikov and Milman [11, 12], which was in turn inspired by [4].
Before presenting our results, we introduce some necessary notation. By a classical characterization due to Borell [2], a measure $\mu$ on $\mathbb{R}^{n}$ is log-concave if and only if there exists a convex function $u:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ such that $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. Throughout this paper, we will always consider $u\in C^{2}(\mathbb{R}^{n})$ such that the Hessian matrix of $u$, denoted by $\nabla^{2}u$, is positive definite.
Let $K\subset\mathbb{R}^{n}$ be a compact and convex set with non-empty interior, such that its boundary $\partial K$ is a manifold of class $C^{2}$ with strictly positive principal curvatures. We shall say that such a set $K$ is of class $C^{2}_{+}$. Adopting the notation introduced by Livshyts [15], the concavity power of $K$ with respect to $\mu$ is the maximal exponent $p(\mu,K)$ such that
$\left.\frac{\,\mathrm{d}^{2}}{\,\mathrm{d}t^{2}}\mu\bigl(K(\rho,t)\bigr)^{p(\mu,K)}\right|_{t=0}\leq 0\quad\text{for every }\rho\in\mathcal{A},$ (2)
where $K(\rho,t)$ denotes a perturbation of $K$ generated by a suitable function $\rho:\partial K\to\mathbb{R}$, and $\mathcal{A}$ is a class of admissible perturbations. We will elaborate more on this point in the next section. By a standard density argument, it can be shown that (dim-BM) holds for $\mu$ if and only if $p(\mu,K)\geq 1/n$ for every origin-symmetric $K$ of class $C^{2}_{+}$, see [15, Section 3.2] for a detailed exposition. Similarly, when studying the validity of (B), we may, by a standard density argument, restrict our attention to compact sets with sufficiently regular boundaries.
Given a non-empty compact and convex set $K\subset\mathbb{R}^{n}$, we recall that its support function is defined as $h_{K}(x)=\sup_{y\in K}\langle x,y\rangle$. Moreover, if $K$ is of class $C^{2}_{+}$, we denote by $\nu_{K}:\partial K\to\mathbb{S}^{n-1}$ its Gauss map. Throughout this note, given a compact and convex set $K$ with non-empty interior and a measure $\mu$ with continuous density with respect to the Lebesgue measure on $K$ (as will always be the case here), we ubiquitously use the symbol $\mathrm{d}\mu$ both for integration on $K$ and on $\partial K$. In the latter case, $\mu$ is understood as its restriction to the (n1)(n-1)-dimensional Hausdorff measure on $\partial K$. We are now ready to state our first main result.
Theorem 1.1.
Let $u\in C^{2}(\mathbb{R}^{n})$ be an even function such that $\nabla^{2}u$ is positive definite, and let $\mu$ be the associated log-concave measure with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. Suppose that for every origin-symmetric set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$,
$\frac{\mu(K)}{\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu}\leq p(\mu,K).$ (3)
Then, for every compact and convex and origin-symmetric set $K$
[…]
In other words, if $\mu$ satisfies (3) for every origin-symmetric $K$ of class $C^{2}_{+}$, $\mu$ satisfies the (B)-conjecture as well.
We say that the even log-concave measure $\mu$ satisfies the strong dimensional Brunn–Minkowski condition if (3) holds for every origin-symmetric set $K$ of class $C^{2}_{+}$. The fact that this condition implies the dimensional Brunn–Minkowski conjecture (dim-BM) follows immediately from an integration by parts, together with the symmetry of the problem. See Corollary 2.2 later.
To state our second result, we introduce some additional notation. Given $u$ and $\mu$ as in Theorem 1.1, for every $\psi\in C^{2}(\mathbb{R}^{n})$ we define the $\mu$-weighted Laplacian as $\mathcal{L}_{\mu}(\psi)=\Delta\psi-\langle\nabla u,\nabla\psi\rangle$. Moreover, we say that a measure $\eta$ is log-concave with respect to $\mu$ if $\eta$ is absolutely continuous with respect to $\mu$ and its corresponding density $\mathrm{d}\eta/\,\mathrm{d}\mu$ is a log-concave function. The notion of hereditary convexity was formally introduced in [5]. We restrict our treatment to the case of log-concave measures: A log-concave measure $\mu$ is hereditarily convex if it is even, and for every even measure $\eta$ which is log-concave with respect to $\mu$ and every even $\psi$ such that $\mathcal{L}_{\mu}(\psi)\in L_{2}(\eta)$, one has
[…]
Here, $||\cdot||_{\rm HS}$ denotes the Hilbert–Schmidt norm of matrices, and $|x|^{2}=\langle x,x\rangle$. As previously mentioned, it is known that the class of hereditarily convex measures includes radially symmetric log-concave measures. It remains an open question of major interest whether every even log-concave measure is hereditarily convex; compare [5, Question 16]. Our second result reads as follows.
Theorem 1.2.
Consider $\mu$ as in Theorem 1.1. If $\mu$ is hereditarily convex, then $\mu$ satisfies the strong dimensional Brunn–Minkowski condition.
We shall prove, in fact, that the strong dimensional Brunn–Minkowski condition follows when $\mu$ is hereditarily convex only with respect to uniform measures on non-empty compact and convex subsets of $\mathbb{R}^{n}$.
To summarize this note, the results presented above show that if $\mu$ is an even-log concave measure as in Theorem 1.1, the implications in [5, (62)] can be refined as follows:
$\mu$ verifies (dim-BM)
$\Longrightarrow$
[…]
$\Longrightarrow$
[…]
2 Preliminaries and proofs of the results
Before starting with the proofs, we introduce and recall some auxiliary notions and results. Consider a compact and convex set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. First, we elaborate on the concavity power defined in (2). There, the perturbations $K(\rho,t)$ are defined as follows: As $K$ is of class $C^{2}_{+}$, it is well-known that for every $\rho\in C^{2}(\partial K)$ there exists $\varepsilon>0$ such that $h_{t}=h_{K}+t\rho(\nu^{-1}_{K})$ are the support functions of compact and convex sets of class $C^{2}_{+}$ for every $|t|<\varepsilon$. We consider the sets $K(\rho,t)$ corresponding to the support functions $h_{t}$. Therefore, if $K$ is of class $C^{2}_{+}$, the family of admissible perturbations is, in fact, $C^{2}(\partial K)$. See also [13, Section 3.2] and, for more details on the theory of convex sets, the monograph by Schneider [16].
Given a set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, we denote its second fundamental form by $\rm II$} and the tangential gradient on $\partial K$ by $\nabla_{\partial K}$. Consider, moreover, a function $u\in C^{2}(\mathbb{R}^{n})$ with strictly positive Hessian matrix and the measure $\mu$ with density $\mathrm{d}\mu(x)=e^{-u(x)}\,\mathrm{d}x$. We write the weighted mean curvature of $\partial K$ as ${\rm H}_{\mu}=\mathrm{tr}({\rm II})-\langle\nabla u,\nu_{K}\rangle$. With a slight abuse of notation, we shall use $\langle\cdot,\cdot\rangle$ to denote also the restriction of the scalar product of $\mathbb{R}^{n}$ to the tangent spaces of $\partial K$. We write $H^{1}(\partial K,\mu)$ for the Sobolev space of functions on $\partial K$ with respect to the measure $\mu$. The following statement, proved in [1, Theorem 1.3], provides an analytic representation for the concavity power $p(\mu,K)$.
Theorem 2.1.
Consider $\mu$ such that it satisfies the assumptions of Theorem 1.1, and a compact and convex origin-symmetric set $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$. Then there exists a unique weak solution $\overline{\rho}\in H^{1}(\partial K,\mu)$ to the equation
[…]
and it satisfies
$p(\mu,K)=\frac{\mu(K)}{\int_{\partial K}\overline{\rho}\,\,\mathrm{d}\mu}.$ (6)
Moreover, $\overline{\rho}$ is even.
For later use, notice that (6) allows one to rewrite (3) as
$\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu\geq\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu.$ (7)
Going back to condition (3), observe that integration by parts yields the identity (see later (16) for a proof)
$\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\mu=n\mu(K)-\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu.$
If we plug this expression in (3), we infer
$p(\mu,K)\geq\frac{1}{n-\frac{1}{\mu(K)}\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}\geq\frac{1}{n},$ (8)
where, for the last inequality, we used that whenever $u$ is even and convex, then $\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu\geq 0$. As anticipated in the introduction, one has the following immediate consequence.
Corollary 2.2.
Consider a log-concave measure $\mu$ as in Theorem 1.1. If $\mu$ satisfies (3) for every compact, convex and origin-symmetric set of class $C^{2}_{+}$, then (dim-BM) holds true.
As a further corollary of (8), considering $tK$ for $t\to 0^{+}$ shows that the infimum of $p(\mu,K)$ on all the possible choices of $K$ cannot give, for $p(\mu,K)$, a lower bound larger than $1/n$. In fact, as mentioned, for example, in [10], an analogous limit argument shows that $1/n\geq\inf_{K}p(\mu,K)$.
We now turn our attention to the PDE (5). For $\rho\in C^{2}(\partial K)$ let us define the operator
$\mathcal{E}(\rho)=-\nabla_{\partial K}\cdot({\rm II}^{-1}\nabla_{\partial K}\rho)+\langle\nabla_{\partial K}u,{\rm II}^{-1}\nabla_{\partial K}{\rho}\rangle-{\rm H}_{\mu}\,\rho+\frac{1}{\mu(K)}\int_{\partial K}\rho\,\,\mathrm{d}\mu.$ (9)
Then, $\mathcal{E}$ is a uniformly elliptic operator on $\partial K$ and $\overline{\rho}$ is a weak solution of $\mathcal{E}(\rho)=1$. We remark that if the support function of $K$ is assumed to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, all the coefficients in (9) are of class at least $C^{1}$. Under this assumption, the function $\overline{\rho}$ is, in fact, a classical solution for (5), as can be seen through classical elliptic regularity results. See, e.g., [8]. Now, observe that integration by parts yields the identity
$\int_{\partial K}\mathcal{E}(\rho)\rho\,\mathrm{d}\mu=\int_{\partial K}\langle{\rm II}^{-1}\nabla_{\partial K}\rho,\nabla_{\partial K}\rho\rangle\,\mathrm{d}\mu-\int_{\partial K}{\rm H}_{\mu}\rho^{2}\,\mathrm{d}\mu+\frac{1}{\mu(K)}\left(\int_{\partial K}\rho\,\mathrm{d}\mu\right)^{2}.$
The right-hand side is non-negative thanks to the weighted Poincaré inequality (compare [12]), which can be obtained by direct differentiation of (1). Therefore,
$\int_{\partial K}\mathcal{E}(\rho)\rho\,\mathrm{d}\mu\geq 0$ (10)
for every $\rho\in C^{2}(\partial K)$. We remark that this inequality holds for less regular functions. Additionally, we remark that the action of $\mathcal{E}$ on the support function, whenever $h_{K}$ is of class $C^{2}$ in $\mathbb{R}^{n}\setminus\{0\}$, is
$\mathcal{E}(h_{K}(\nu_{K}))=1+\langle\nabla u,x\rangle-\frac{1}{\mu(K)}\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu.$ (11)
This equality can be found in [1, (67)]. It follows from direct computations, which we omit for the sake of brevity.
Next, we recall a local formulation of the (B)-conjecture. Differentiating twice $\mu(e^{t}K)$ for $\mu$ and $K$ as above (see, for example, [6, (10)], which in turn generalizes [4, (4)]), shows that the log-concavity of the function $t\mapsto\mu(e^{t}K)$ at $t=0$ is equivalent to
$\frac{\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\mu}{\mu(K)}-\left(\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}\right)^{2}\leq\frac{\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\mu}{\mu(K)}+\frac{\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\mu}{\mu(K)}.$ (12)
Thus, (B) is equivalent to the validity of (12) for every compact and origin-symmetric convex set $K\subset\mathbb{R}^{n}$ with non-empty interior. For the $\mu$-weighted Laplacian $\mathcal{L}_{\mu}$, we recall the following useful integration by parts identity, which holds for any two functions $\varphi,\psi\in C^{2}(K)$:
$\int_{K}\varphi\,\mathcal{L}_{\mu}(\psi)\,\mathrm{d}\mu=-\int_{K}\langle\nabla\varphi,\nabla\psi\rangle\,\mathrm{d}\mu+\int_{\partial K}\varphi\langle\nabla\psi,\nu_{K}\rangle\,\mathrm{d}\mu.$ (13)
Finally, we will make use of the following Reilly formula, proved in [11, (2.3)] in the more general context of weighted Riemannian manifolds (compare also [5, (21)]). Consider a function $\psi\in C^{2}(\mathbb{R}^{n})$ and its normal derivative $\rho=\langle\nabla\psi,\nu_{K}\rangle$. Then
$\int_{K}(\mathcal{L}_{\mu}\psi)^{2}\,\mathrm{d}\mu$ $=\int_{K}\|\nabla^{2}\psi\|_{{\rm HS}}^{2}+\langle(\nabla^{2}u)\nabla\psi,\nabla\psi\rangle\,\mathrm{d}\mu$
$+\int_{\partial K}{\rm H}_{\mu}\rho^{2}+\langle{\rm II}\nabla_{\partial K}\psi,\nabla_{\partial K}\psi\rangle-2\langle\nabla_{\partial K}\psi,\nabla_{\partial K}\rho\rangle\,\mathrm{d}\mu.$ (14)
We are now ready to prove our results.
Proof of Theorem 1.1.
We begin by following the argument in [1, Proposition 5.3]. Let $K\subset\mathbb{R}^{n}$ be a compact and convex origin-symmetric set and let $\overline{\rho}$ be given by (5). We may assume by approximation that, in addition, the support function of $K$ is of class $C^{3}$ with strictly positive principal curvatures and, thus, $\overline{\rho}\in C^{2}(\partial K)$. Moreover, denote by $\eta$ the measure with density $\mathrm{d}\eta=\frac{1}{\mu(K)}\mathbbm{1}_{K}\,\mathrm{d}\mu$, where $\mathbbm{1}_{K}$ is the characteristic function of $K$ such that $\mathbbm{1}_{K}(x)=1$ if $x\in K$ and 0 otherwise. Suppose that (3) holds, we shall show that (12) is also satisfied. From (11) and since $\mathcal{E}(\overline{\rho})=1$ we infer
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\int_{\partial K}(h_{K}(\nu_{K})-\overline{\rho})\mathcal{E}(\overline{\rho})\,\mathrm{d}\eta=\int_{\partial K}\mathcal{E}(h_{K}(\nu_{K})-\overline{\rho})\overline{\rho}\,\mathrm{d}\eta$
$=\int_{\partial K}\overline{\rho}\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta.$
Then, writing $f(x)=\overline{\rho}(x)-h_{K}(\nu_{K}(x))$,
$\int_{\partial K}h_{K}(\nu_{K})-\overline{\rho}\,\mathrm{d}\eta$ $=\underbrace{\int_{\partial K}f\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}f\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm I})}$
$+\underbrace{\int_{\partial K}h_{K}(\nu_{K})\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta}_{({\rm I\!I})}.$ (15)
We begin by computing the term $({\rm I})$. Notice that by (11) and (5) we have
$\mathcal{E}(f)=\mathcal{E}(\overline{\rho}-h_{K}(\nu_{K}))=-\langle\nabla u,x\rangle+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta.$
Thus,
$({\rm I})=-\int_{\partial K}f\mathcal{E}(f)\,\mathrm{d}\eta\leq 0\,$,
where the last inequality is (10). We proceed with the term $({\rm I\!I})$. Notice that since for every $x\in\partial K$ we have that $h_{K}(\nu_{K}(x))=\langle x,\nu_{K}(x)\rangle=\langle\nabla(|x|^{2}/2),\nu_{K}(x)\rangle$, an application of (13) implies
$\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta=\int_{K}\mathcal{L}_{\mu}\left(\frac{|x|^{2}}{2}\right)\,\mathrm{d}\eta=n-\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta.$ (16)
Moreover, integrating again by parts shows
$\int_{\partial K}\langle\nabla u,x\rangle h_{K}(\nu_{K})\,\mathrm{d}\eta$ $=\int_{K}\mathcal{L}_{\mu}\left(\frac{|x|^{2}}{2}\right)\langle\nabla u,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla\langle\nabla u,x\rangle,x\big\rangle\,\mathrm{d}\eta$
$=n\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\eta+\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta.$
Therefore, we infer
$({\rm I\!I})=\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta-\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\eta+\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta\right)^{2}.$ (17)
If the strong dimensional Brunn–Minkowski condition holds, by (7) we have that $({\rm I})+({\rm I\!I})\geq 0$ and, since $({\rm I})\leq 0$, we conclude that $({\rm I\!I})\geq 0$. That is,
$\int_{K}\langle(\nabla^{2}u)x,x\rangle\,\mathrm{d}\eta+\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta\geq\int_{K}\langle\nabla u,x\rangle^{2}\,\mathrm{d}\eta-\left(\int_{K}\langle\nabla u,x\rangle\,\mathrm{d}\eta\right)^{2},$
which is precisely (12). Since $K$ was arbitrary we infer that (B) holds, concluding the proof.
Proof of Theorem 1.2.
For this proof, we apply Hörmander’s $L_{2}$ method, following the approach introduced in [11, 12], which can be consulted for further details. See also the discussion in [5, Section 2]. Given $K\subset\mathbb{R}^{n}$ of class $C^{2}_{+}$, we shall prove that (3) is satisfied. Consider the function $\overline{\rho}$ solving (5) for the previous choice of $K$, and let $\psi$ be the unique solution to the Neumann problem
$\begin{cases}\mathcal{L}_{\mu}\psi=\frac{1}{\mu(K)}\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu&\text{in }K,\\[2.0pt] \displaystyle\langle\nabla\psi,\nu_{K}\rangle=\overline{\rho}&\text{on }\partial K.\end{cases}$ (18)
Such solution exists provided that the integral (with respect to $\mu$) of the interior data on $K$ is the same as the integral of the boundary data on $\partial K$, which is indeed the case here. By approximation, we can assume $h_{K}$ to be of class $C^{3}$ in $\mathbb{R}^{n}\setminus\{0\}$, so that $\overline{\rho}\in C^{2}(\partial K)$. Compare the discussion after (9). Thus, standard results in elliptic regularity theory (see, for example, [11, Theorem 2.5]) yield that $\psi\in C^{2}(K)$. Consider the normalized measure $\eta$ as in the previous proof. The Cauchy–Schwarz inequality for positive quadratic forms implies that
$0\leq\langle{\rm II}\nabla_{\partial K}\psi,\nabla_{\partial K}\psi\rangle-2\langle\nabla_{\partial_{K}}\psi,\nabla_{\partial K}\overline{\rho}\rangle+\langle{\rm II}^{-1}\nabla_{\partial K}\overline{\rho},\nabla_{\partial K}\overline{\rho}\rangle.$
Applying this to the Reilly formula (2), for $\mu$, $\psi$ and dividing by $\mu(K)$ implies
$\int_{\partial K}\langle{\rm II}^{-1}\nabla_{\partial K}\overline{\rho},\nabla_{\partial K}\overline{\rho}\rangle\,\mathrm{d}\eta-\int_{\partial K}{\rm H}_{\mu}\overline{\rho}^{2}\,\mathrm{d}\eta+\left(\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\right)^{2}\geq\int_{K}\|\nabla\psi\|^{2}_{{\rm HS}}+\langle(\nabla u)^{2}\nabla\psi,\nabla\psi\rangle\,\mathrm{d}\eta,$ (19)
Furthermore, since by (5) $\overline{\rho}$ satisfies $\mathcal{E}(\overline{\rho})=1$, multiplying with $\overline{\rho}$ and integrating yields
$\int_{\partial K}\langle{\rm II}^{-1}\nabla_{\partial K}\overline{\rho},\nabla_{\partial K}\overline{\rho}\rangle\,\mathrm{d}\eta-\int_{\partial K}{\rm H}_{\mu}\rho^{2}\,\mathrm{d}\eta+\left(\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\right)^{2}=\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta.$ (20)
Notice that this implies that $\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\geq 0$. Furthermore, by [1, Corollary 3.4] equality in the corresponding Poincaré inequality holds if and only if the function is identically zero. Consequently, we infer that $\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta>0$, as the alternative $\overline{\rho}\equiv 0$ would contradict (5). Now, notice that
$\int_{K}\mathcal{L}_{\mu}\psi\,\mathrm{d}\eta=\int_{K}\frac{1}{\mu(K)}\int_{\partial K}\overline{\rho}\,\mathrm{d}\mu\,\mathrm{d}\eta=\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta,$
by definition of the measure $\eta$. Also, by Theorem 2.1, $\overline{\rho}$ is even and thus so is $\psi$. Furthermore, since $\int_{K}\mathcal{L}_{\mu}(\frac{x^{2}}{2})\,\mathrm{d}\eta=\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta$ by (16), we conclude by the hereditary convexity assumption (4) that
$\int_{K}\|\nabla\psi\|^{2}_{{\rm HS}}+\langle(\nabla u)^{2}\nabla\psi,\nabla\psi\rangle\,\mathrm{d}\nu\geq\frac{\left(\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\right)^{2}}{\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta},$
which, in combination with (19), (20), and the fact that $\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta>0$ implies
$\int_{\partial K}\overline{\rho}\,\mathrm{d}\eta\leq\int_{\partial K}h_{K}(\nu_{K})\,\mathrm{d}\eta.$
Thus, (3) holds for an arbitrary compact, convex and origin-symmetric set $K$ and the strong dimensional Brunn–Minkowski condition holds for the measure $\mu$, concluding the proof.
Acknowledgments.
The first-named author was supported by the Austrian Science Fund (FWF): 10.55776/P36344N. The second-named author was supported, in part, by the Austrian Science Fund (FWF): 10.55776/PAT3787224.